# Qudit-native measurement protocol for dynamical correlations using Hadamard tests

### v1

@@ III.2.1 Implementation of the controlled unitaries for spin observables
In order to implement our protocol, we need to calculate the explicit form of the unitary operators in the decomposition of Eq. (3) for the relevant spin-1 operators. With respect to the eigenbasis of the $\hat{S}_{z}$ operator, we find in matrix representation
@@ III.2.2 Performance benchmark in the presence of shot noise
In an experimental setting, the expectation values in Eq. (10) as well as the outcome of the Hadamard tests are subject to errors caused by finite measurement statistics. We simulate these shot-noise errors by sampling the observables from the ideal distribution given by the total wave function of the system–ancilla register after the execution of the Hadamard-test circuit. To this end, we assign a total budget of shots for measuring the four correlation functions on the right hand side of Eq. (5) as well as the two expectation values forming the disconnected part of the anti-commutator.
Figure 2 compares the exact value of $C_{12}^{\pm}(0,t)$ (blue solid line) to the one obtained from sampling with finite statistics (blue circles). The shot noise is low enough to allow for an accurate measurement of the two-time anti-commutator (commutator) already for a moderate number of $250$ ($2000$) shots for each expectation value, resulting in a total number of $1500$ ($8000$) shots per point.
@@ III.3 Comparison to protocols based on linear response
Figure 3: Bias–variance tradeoff in the linear-response measurement of dynamical correlations. With increasing perturbation strength $\lambda$, the linear-response protocol incurs systematic errors in the two-time anti-commutator (a) and commutator (b) due to nonlinear effects. Concomitantly, the statistical error (standard deviation of a single measurement) of the anti-commutator (c) and commutator (d) decreases, requiring a tradeoff between accuracy and signal-to-noise ratio. By contrast, the Hadamard-test protocol exhibits consistently lower statistical errors and does not introduce intrinsic systematic sources of error.
In this subsection, we compare our Hadamard-test protocol with other protocols for measuring dynamical correlation functions. Many such protocols are specific to certain platforms or observables. To keep our analysis general, we focus the following discussion on protocols based on linear response theory [18, 19]. In fact, a large class of well-established protocols for measuring dynamical correlations, including non-invasive (or weak) measurements [43], can be phrased in an implementation-independent way using this framework [25].
@@
As well known in linear response theory, the response of the system to a weak (Hermitian) perturbation gives access to the two-time commutator according to Kubo’s formula [18, 19]. More recently, it has been proposed to measure the two-time anti-commutator by probing the linear response to a non-Hermitian perturbation [56, 57, 25]. Here, we employ such linear-response protocols to extract the response function directly in the time domain and benchmark the result against our Hadamard-test protocol.
Specifically, to measure the two-time (anti-)commutator $C_{ij}^{\pm}(t_{1},t_{2})$, the system is first evolved under the unperturbed Hamiltonian $\smash{\hat{H}_{0}}$ to the time $t_{1}$. Then, a Hermitian (anti-Hermitian) perturbation of strength $0<\lambda\ll 1$ by the operator $\hat{S}_{j}^{z}$ is applied in the form of a short rectangular pulse of duration $\Delta t$, during which the system evolves under the perturbed Hamiltonian $\hat{H}_{\lambda}=\hat{H}_{0}-\lambda\hbar J_{xy}\hat{S}_{j}^{z}$ (in the non-Hermitian case, we have $\lambda\to i\lambda$). After releasing the perturbation, the system is evolved under the unperturbed Hamiltonian $\smash{\hat{H}_{0}}$ up to time $t_{2}$ and the expectation value of the observable $\hat{S}_{i}^{z}$ is measured, yielding the two-time (anti-)commutator via
$C^{\pm}_{ij,\,\mathrm{LR}}(t_{1},t_{2})$ $=\frac{1}{\lambda J_{xy}\Delta t}\big{(}\big{\langle}\hat{S}^{z}_{i}(t_{2})\big{\rangle}_{\pm}-\big{\langle}\hat{S}^{z}_{i}(t_{2})\big{\rangle}\big{)}.$ (10)
@@
III.3.2 Finite statistics and bias–variance tradeoff
As depicted in Fig. 2, the linear-response protocol accurately reproduces the time trace of the dynamical correlations for a sufficiently weak perturbation and infinite measurement statistics (red solid line). An advantage of our Hadamard-test protocol over the linear-response method becomes apparent if one takes statistical errors due to shot noise into account. The red circles show the sampled values of the anti-commutator (commutator) using $750$ ($6000$) shots for each expectation value in Eq. 10, corresponding to a total number of $1500$ ($12000$) shots per point, which is comparable to the total budget used for the Hadamard-test protocol  [^1] In the non-Hermitian case, we reduce the number of shots according to the decrease of the norm of the state.. One can see that the Hadamard-test protocol yields a lower statistical variance and therefore a higher signal-to-noise ratio.
@@ III.3.2 Finite statistics and bias–variance tradeoff
The linear-response protocol generally requires a tradeoff between accuracy and signal-to-noise ratio. On the one hand, the perturbation should be sufficiently weak such that the response is in the linear regime where nonlinear contributions to the response are negligible. On the other hand, it should be sufficiently strong such that the signal is resolvable against the background noise. As can be seen in Fig. 3, the error grows with increasing $\lambda$, while the variance shrinks, and vice versa. This behavior can be viewed as a manifestation of the bias–variance tradeoff, well known in statistics and machine learning. Importantly, there is no such tradeoff in the case of our Hadamard-test protocol. Furthermore, in the investigated quench scenario, the Hadamard-test protocol outperforms the linear-response protocol both in terms of accuracy and signal-to-noise ratio.
@@ III.3.3 Experimental resources
For the commutator, the linear-response protocol requires a standard Hermitian perturbation, which may be either single-body, as in the case of spin–spin dynamical correlations, or involve few-body interactions. Importantly, the protocol typically admits ancilla-free realizations, where the perturbation is exerted directly on the system, e.g., in the form of a rapid change of the couplings in the Hamiltonian or as a unitary gate in a quantum circuit. By contrast, the protocol based on Hadamard tests requires an ancillary qubit and two entangling operations between this qubit and the system. Thus, depending on the device and scenario at hand, a decision has to be made between additional errors from shot noise and bias (for the linear-response protocol, see Fig. 3), and potential errors from faulty gates (for the Hadamard-test protocol). One way of circumventing the need of a physical ancilla for the Hadamard-test is to use qudits with a number of controllable levels equal to double the local Hilbert space dimension of the system we want to simulate (e.g., $d=6$ in the example above).
In our benchmarks of the anti-commutator, we have considered an implementation-agnostic “ideal” non-Hermitian linear-response measurement, where a non-Hermitian perturbation directly acts on the system degrees of freedom. However, physical realizations of this scenario typically incur additional experimental overhead, e.g., entangling the system with an ancilla and performing postselected measurements on the ancilla state [25]. Therefore, engineering an (approximate) effective non-Hermitian Hamiltonian requires similar resources as the Hadamard-test protocol.
@@ IV Conclusion
In this work, we have proposed an algorithm for measuring dynamical correlation functions on qudit quantum hardware. Key to the scheme is a decomposition of general (multi-)qudit observables into unitaries, which can be exploited to represent dynamical correlations as a sum of quantities accessible via Hadamard tests. We have numerically benchmarked the performance of this protocol with finite measurement statistics at the example of a spin-1 XXZ model after a quench. Notably, in this scenario the Hadamard-test protocol outperforms established protocols based on linear response theory both in terms of accuracy and number of shots needed to distinguish the signal from noise.
The Hadamard-test protocol for dynamical correlation functions can be extended in a straightforward way to measuring higher-order temporal correlations of the form $\langle A_{1}(t_{1})\cdots A_{n}(t_{n})\rangle$. Such a generalization of our protocol would only require the implementation of deeper circuits and a higher number of controlled operations, while the structure of the circuit and the usage of the decomposition of Hermitian operators into a sum of unitaries remain the same.
Finally, we emphasize that our Hadamard-test protocol for measuring dynamical correlations in qudit quantum systems can be used as a subroutine in existing algorithms with wide-ranged applications, ranging from thermalization in many-body systems [24, 25, 59, 52] to the implementation of variational optimization protocols for the simulation of equilibrium and non-equilibrium properties of many-body systems [39, 35, 40]. Moreover, the protocol is designed such that it can be implemented on presently available quantum simulation platforms, including trapped ions, cold atoms, and superconducting qubits.
@@ V Acknowledgements
ICFO group acknowledges support from: ERC AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033,  CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB-I00 project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA MAQS PCI2019-111828-2);  QUANTERA DYNAMITE PCI2022-132919 (QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733), Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation, and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-1-0013); EU Quantum Flagship (PASQuanS2.1, 101113690); EU Horizon 2020 FET-OPEN OPTOlogic (Grant No 899794); EU Horizon Europe Program (Grant Agreement 101080086 — NeQST), ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID 100010434): CF/BQ/PR23/11980043.
This project has received funding from: the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST, the European Union under NextGenerationEU PRIN 2022 Prot. n. 2022ATM8FY (CUP: E53D23002240006), the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A), the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.4 - Call for tender No. 1031 of 17/06/2022 of Italian Ministry for University and Research funded by the European Union – NextGenerationEU (proj. nr. CN_00000013), and Project DYNAMITE QUANTERA2_00056 funded by the Ministry of University and Research through the ERANET COFUND QuantERA II – 2021 call and co-funded by the European Union (H2020, GA No 101017733). This work has benefited from Q@TN, the joint lab between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. We acknowledge support by Provincia Autonoma di Trento. P.P.P. acknowledges also support from the “Secretaria d’Universitats i Recerca del Departament de Recerca i Universitats de la Generalitat de Catalunya” under grant 2024 FI-3 00390, as well as the European Social Fund Plus.
@@ Appendix A Outcome of the quantum circuit implementing the Hadamard-test protocol
with $\alpha\in\{0,\frac{\pi}{2}\}$, depending on whether we want to measure the unequal-time anti-commutator or commutator. After flipping the ancilla, evolving the state up to time $t_{1}$, and applying the first controlled operation $V_{A}$, we obtain the intermediate state
$\frac{1}{\sqrt{2}}e^{i\alpha}|0\rangle\otimes U(t_{1})|\psi_{0}\rangle+\frac{1}{\sqrt{2}}|1\rangle\otimes V_{A}U(t_{1})|\psi_{0}\rangle.$ (14)
Next, we apply the second ancilla flip, evolve the register up to time $t_{2}$, and apply the second controlled operation $V_{B}$, yielding
$\begin{split}&\frac{1}{\sqrt{2}}|0\rangle\otimes U(t_{2}-t_{1})V_{A}U(t_{1})|\psi_{0}\rangle\\ +&\frac{1}{\sqrt{2}}e^{i\alpha}|1\rangle\otimes V_{B}U(t_{2})|\psi_{0}\rangle.\end{split}$ (15)
@@
$\begin{split}&|\Psi_{V_{A}V_{B}}^{\alpha}(t_{1},t_{2})\rangle=\\ &|0\rangle\otimes\left[\frac{1}{2}U(t_{2}-t_{1})V_{1}U(t_{1})|\psi_{0}\rangle+\frac{1}{2}e^{i\alpha}V_{2}U(t_{2})|\psi_{0}\rangle\right]\\ +&|1\rangle\otimes\left[\frac{1}{2}U(t_{2}-t_{1})V_{1}U(t_{1})|\psi_{0}\rangle-\frac{1}{2}e^{i\alpha}V_{2}U(t_{2})|\psi_{0}\rangle\right].\end{split}$ (16)
@@
$4\left|\left\langle 0\middle|\mathopen{}\Psi_{V_{A}V_{B}}^{\alpha}(t_{1},t_{2})\right\rangle\right|^{2}-2\\ =e^{i\alpha}\langle\psi_{0}|U^{\dagger}(t_{2})V^{\dagger}_{2}U(t_{2}-t_{1})V_{1}U(t_{1})|\psi_{0}\rangle+\text{H.c.}$ (17)
@@
By setting $\alpha=0$ or $\alpha=\frac{\pi}{2}$, we get the real or the imaginary part of the dynamical correlation function of $V_{A}$ and $V^{\dagger}_{B}$, respectively. This proves Eq. (6).
Appendix B Variance of the Hadamard-test
We proceed with the proof of Eq. (7). Due to Eq. (6), the variance of the dynamical correlation of interest satisfies $\operatorname{Var}[C^{\pm}_{XY}]=16\operatorname{Var}[\hat{\Pi}_{0}]$, where
$\mathrm{Var}[\hat{\Pi}_{0}]=\langle\hat{\Pi}_{0}^{2}\rangle-\langle\hat{\Pi}_{0}\rangle^{2}.$ (18)
Here, $\hat{\Pi}_{0}=|0\rangle\langle 0|\otimes\mathds{1}$ denotes the projector onto the ancilla state $|0\rangle$ and expectation values are calculated in the state $|\Psi_{V_{1}V_{2}}^{\alpha}(t_{1},t_{2})\rangle$. For the projector $\hat{\Pi}_{0}$, the simple identity $\hat{\Pi}_{0}^{2}=\hat{\Pi}_{0}$ holds, and thus
$\mathrm{Var}[\hat{\Pi}_{0}]=\langle\hat{\Pi}_{0}\rangle(1-\langle\hat{\Pi}_{0}\rangle).$ (19)

### v2

@@ III.2.1 Implementation of the controlled unitaries for spin observables
In order to implement our protocol, we need to calculate the explicit form of the unitary operators in the decomposition of Eq. 3. While for a general qudit observable $\hat{X}$ the structure of the resulting unitary $\hat{W}_{X}$ may be complicated, the situation simplifies for spin observables. The relevant operator for our case study is the single-spin $\hat{S}_{z}$ operator. In the general spin-$S$ case, the matrix elements with respect to the eigenbasis of the $\hat{S}_{z}$ operator take the simple form
$\big{\langle}m\,\big{|}\,\mathopen{}\hat{W}_{S_{z}}\,\big{|}\,\mathopen{}m^{\prime}\big{\rangle}=\delta_{m,m^{\prime}}\left(\frac{m}{S}+i\sqrt{1-\frac{m^{2}}{S^{2}}}\right)\,,$ (9)
where $m,m^{\prime}\in\{S,\dots,-S\}$. In particular, for $S=1$, we find in matrix representation
@@ III.2.2 Performance benchmark in the presence of shot noise
In an experimental setting, the expectation values in Eq. (11) as well as the outcome of the Hadamard tests are subject to errors caused by finite measurement statistics. We simulate these shot-noise errors by sampling the observables from the ideal distribution given by the total wave function of the system–ancilla register after the execution of the Hadamard-test circuit. To this end, we assign a total budget of shots for measuring the four probabilities on the right-hand side of Eq. 6 as well as the two expectation values forming the disconnected part of the anti-commutator.
Figure 2 compares the exact value of $C_{12}^{\pm}(0,t)$ (blue solid line) to the one obtained from sampling with finite statistics (blue circles). The shot noise is low enough to allow for an accurate measurement of the two-time anti-commutator (commutator) already for a moderate number of $250$ ($2000$) shots for each expectation value, resulting in a total of $1500$ ($8000$) shots per point.
@@ III.3 Comparison to protocols based on linear response
Figure 3: Bias–variance tradeoff in the linear-response measurement of dynamical correlations. With increasing perturbation strength $\lambda$, the linear-response protocol incurs systematic errors in the two-time (a) anti-commutator and (b) commutator due to nonlinear effects. Concomitantly, the statistical error (standard deviation of a single measurement) of the (c) anti-commutator and (d) commutator decreases, requiring a trade-off between accuracy and signal-to-noise ratio. By contrast, the Hadamard-test protocol exhibits consistently lower statistical errors and does not introduce intrinsic systematic sources of error.
In this section, we compare our Hadamard-test protocol with other protocols for measuring dynamical correlation functions. Many such protocols are specific to certain platforms or observables. To keep our analysis general, we focus the following discussion on protocols based on linear response theory [18, 19]. In fact, a large class of well-established protocols for measuring dynamical correlations, including non-invasive (or weak) measurements [44], can be phrased in an implementation-independent way using this framework [25].
@@
As well known in linear response theory, the response of the system to a weak (Hermitian) perturbation gives access to the two-time commutator according to Kubo’s formula [18, 19]. More recently, measuring the two-time anti-commutator by probing the linear response to a non-Hermitian perturbation was proposed [62, 63, 25]. Here, we employ such linear-response protocols to extract the response function directly in the time domain and benchmark the result against our Hadamard-test protocol.
Specifically, to measure the two-time (anti-)commutator $C_{ij}^{\pm}(t_{1},t_{2})$, the system is first evolved under the unperturbed Hamiltonian $\smash{\hat{H}_{0}}$ to the time $t_{1}$. Then, a Hermitian (anti-Hermitian) perturbation of strength $0<\lambda\ll 1$ by the operator $\hat{S}_{j}^{z}$ is applied in the form of a short rectangular pulse of duration $\Delta t$, during which the system evolves under the perturbed Hamiltonian $\hat{H}_{\lambda}=\hat{H}_{0}-\lambda\hbar J_{xy}\hat{S}_{j}^{z}$ (in the non-Hermitian case, we have $\lambda\to i\lambda$). After the perturbation is released, the system is evolved under the unperturbed Hamiltonian $\smash{\hat{H}_{0}}$ up to time $t_{2}$, and the expectation value of the observable $\hat{S}_{i}^{z}$ is measured, yielding the two-time (anti-)commutator via
$C^{\pm}_{ij,\,\mathrm{LR}}(t_{1},t_{2})$ $=\frac{1}{\lambda J_{xy}\Delta t}\big{[}\big{\langle}\hat{S}^{z}_{i}(t_{2})\big{\rangle}_{\pm}-\big{\langle}\hat{S}^{z}_{i}(t_{2})\big{\rangle}\big{]}\,.$ (11)
@@
III.3.2 Finite statistics and bias–variance trade-off
As depicted in Fig. 2, the linear-response protocol accurately reproduces the time trace of the dynamical correlations for a sufficiently weak perturbation and infinite measurement statistics (red solid line). An advantage of our Hadamard-test protocol over the linear-response method becomes apparent if one takes statistical errors due to shot noise into account. The red circles show the sampled values of the anti-commutator (commutator) using $750$ ($6000$) shots for each expectation value in Eq. 11, corresponding to a total of $1500$ ($12000$) shots per point, which is comparable to the total budget used for the Hadamard-test protocol  [^1] In the non-Hermitian case, we reduce the number of shots according to the decrease of the norm of the state.. One can see that the Hadamard-test protocol yields a lower statistical variance and therefore a higher signal-to-noise ratio.
@@ III.3.2 Finite statistics and bias–variance trade-off
The linear-response protocol generally requires a trade-off between accuracy and signal-to-noise ratio. On the one hand, the perturbation should be sufficiently weak that the response is in the linear regime, where nonlinear contributions to the response are negligible. On the other hand, it should be sufficiently strong that the signal is resolvable against the background noise. As can be seen in Fig. 3, the error grows with increasing $\lambda$, while the variance shrinks and vice versa. This behavior can be viewed as a manifestation of the bias–variance tradeo-ff, which is well-known in statistics and machine learning. Importantly, there is no such trade-off in the case of our Hadamard-test protocol. Furthermore, in the investigated quench scenario, the Hadamard-test protocol outperforms the linear-response protocol both in terms of accuracy and signal-to-noise ratio.
@@ III.3.3 Experimental resources
For the commutator, the linear-response protocol requires a standard Hermitian perturbation, which either may be single-body, as in the case of spin–spin dynamical correlations, or may involve few-body interactions. Importantly, the protocol typically admits ancilla-free realizations, where the perturbation is exerted directly on the system, e.g., in the form of a rapid change in the couplings in the Hamiltonian or as a unitary gate in a quantum circuit. By contrast, the protocol based on Hadamard tests requires an ancillary qubit and two entangling operations between this qubit and the system. Thus, depending on the device and scenario at hand, a decision has to be made between additional errors from shot noise and bias (for the linear-response protocol; see Fig. 3) and potential errors from faulty gates (for the Hadamard-test protocol). One way of circumventing the need for a physical ancilla for the Hadamard test is to use qudits with the number of controllable levels equal to double the local Hilbert space dimension of the system we want to simulate (e.g., $d=6$ in the example above).
In our benchmarks of the anti-commutator, we have considered an implementation-agnostic “ideal” non-Hermitian linear-response measurement, where a non-Hermitian perturbation directly acts on the system degrees of freedom. However, physical realizations of this scenario typically incur additional experimental overhead, e.g., entangling the system with an ancilla and performing postselected measurements on the ancilla state [25]. Therefore, engineering an (approximate) effective non-Hermitian Hamiltonian requires resources similar to those of the Hadamard-test protocol.
Finally, we turn our attention to potential near-term experimental realizations of our Hadamard-test protocol for dynamical correlations of general observables. For systems with a large Hilbert space, implementing a general multi-qudit observable as a sum of unitaries presents significant practical challenges. While the decomposition in Eq. 3 applies to any observable, the resulting unitary may have a complicated structure (unlike for spin observables; see Section III.2.1). In this case, the gate complexity may be reduced by resorting to approximate decompositions, e.g., based on the LCU procedure [54], where one can trade off lower circuit depth for a more accurate approximation. While in typical applications the circuit depth is dominated by the number of gates required to implement the system’s time evolution, the added gate complexity due to the (approximate) implementation of the controlled unitaries may be of concern on noisy intermediate-scale quantum (NISQ) devices: In the presence of circuit noise, the maximum circuit depth is restricted by the finite coherence time. This may require lowering the approximation accuracy of the controlled unitaries, introducing systematic errors, and it may impose limits on the times $t_{1}$ and $t_{2}$ of the dynamical correlations that can reliably be measured using the Hadamard test protocol. In situations where this effect is identified as a limiting factor, the linear-response protocol may provide a viable alternative since the implementation of the perturbation is typically favorable in terms of gate complexity.
@@ IV Conclusion
In this work, we proposed an algorithm for measuring dynamical correlation functions on qudit quantum hardware. Key to the scheme is the decomposition of general (multi-)qudit observables into unitaries, which can be exploited to represent dynamical correlations as the sum of quantities accessible via Hadamard tests. We numerically benchmarked the performance of this protocol with finite measurement statistics using the example of a spin-1 XXZ model after a quench. Notably, in this scenario the Hadamard-test protocol outperforms established protocols based on linear response theory in terms of both accuracy and the number of shots needed to distinguish the signal from noise.
The Hadamard-test protocol for dynamical correlation functions can be extended in a straightforward way to measuring higher-order temporal correlations of the form $\langle A_{1}(t_{1})\cdots A_{n}(t_{n})\rangle$. Such a generalization of our protocol would require only the implementation of deeper circuits and a higher number of controlled operations, while the structure of the circuit and the usage of the decomposition of Hermitian operators into a sum of unitaries remain the same.
Finally, we emphasize that our Hadamard-test protocol for measuring dynamical correlations in qudit quantum systems can be used as a subroutine in existing algorithms with wide-ranging applications, ranging from thermalization in many-body systems [24, 25, 65, 53] to the implementation of variational optimization protocols for the simulation of equilibrium and non-equilibrium properties of many-body systems [40, 36, 41]. Moreover, the protocol is designed such that it can be implemented on presently available quantum simulation platforms, including trapped ions, cold atoms, and superconducting qubits.
@@ V Acknowledgements
The ICFO group acknowledges support from: ERC AdG NOQIA; MCIN/AEI [PGC2018-0910.13039/501100011033,  CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, the Plan National STAMEENA PID2022-139099NB-I00 project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI]; QUANTERA MAQS PCI2019-111828-2);  QUANTERA DYNAMITE PCI2022-132919 (QuantERA II Programme co-funded by the European Union’s Horizon 2020 program under Grant Agreement No. 101017733), Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and the European Union through the Recovery, Transformation, and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-1-0013); EU Quantum Flagship (PASQuanS2.1, 101113690); EU Horizon 2020 FET-OPEN OPTOlogic (Grant No. 899794); the EU Horizon Europe Program (Grant Agreement No. 101080086 — NeQST), ICFO Internal “QuantumGaudi” project; the European Union’s Horizon 2020 program under Marie Sklodowska-Curie Grant Agreement No. 847648; and “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID No. 100010434): CF/BQ/PR23/11980043.
This project received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101080086 NeQST, the European Union under NextGenerationEU PRIN 2022 Prot. n. 2022ATM8FY (CUP: E53D23002240006), the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant No. R20PEX7Y3A), the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.4 - Call for tender No. 1031 of the Italian Ministry for University and Research funded by the European Union – NextGenerationEU (Project No. CN_00000013), and Project DYNAMITE QUANTERA2_00056 funded by the Ministry of University and Research through the ERANET COFUND QuantERA II – 2021 call and co-funded by the European Union (H2020, Grant Agreement No. 101017733). This work benefited from Q@TN, the joint laboratory between the University of Trento, Fondazione Bruno Kessler (FBK), National Institute for Nuclear Physics (INFN), and National Research Council (CNR). We acknowledge support from Provincia Autonoma di Trento. P.P.P. also acknowledges support from the “Secretaria d’Universitats i Recerca del Departament de Recerca i Universitats de la Generalitat de Catalunya” under Grant No. 2024 FI-3 00390, as well as the European Social Fund Plus.
@@
Data Availability
The data set used in this paper is available on Zenodo [66].
@@ Appendix A Outcome of the quantum circuit implementing the Hadamard-test protocol
with $\alpha\in\{0,\frac{\pi}{2}\}$, depending on whether we want to measure the unequal-time anti-commutator or commutator. After flipping the ancilla, evolving the state up to time $t_{1}$, and applying the first controlled operation $\hat{V}_{A}$, we obtain the intermediate state
$\frac{1}{\sqrt{2}}e^{i\alpha}|0\rangle\otimes\hat{U}(t_{1})|\psi_{0}\rangle+\frac{1}{\sqrt{2}}|1\rangle\otimes\hat{V}_{A}\hat{U}(t_{1})|\psi_{0}\rangle\,.$ (15)
Next, we apply the second ancilla flip, evolve the register up to time $t_{2}$, and apply the second controlled operation $\hat{V}_{B}$, yielding
$\begin{split}\frac{1}{\sqrt{2}}|0\rangle&\otimes\hat{U}(t_{2}-t_{1})\hat{V}_{A}\hat{U}(t_{1})|\psi_{0}\rangle\\ +\frac{1}{\sqrt{2}}e^{i\alpha}|1\rangle&\otimes\hat{V}_{B}\hat{U}(t_{2})|\psi_{0}\rangle\,.\end{split}$ (16)
@@
$\begin{split}&|\Psi_{V_{A}V_{B}}^{\alpha}(t_{1},t_{2})\rangle\\ =&|0\rangle\otimes\left[\frac{1}{2}\hat{U}(t_{2}-t_{1})\hat{V}_{A}\hat{U}(t_{1})|\psi_{0}\rangle+\frac{1}{2}e^{i\alpha}\hat{V}_{B}\hat{U}(t_{2})|\psi_{0}\rangle\right]\\ +&|1\rangle\otimes\left[\frac{1}{2}\hat{U}(t_{2}-t_{1})\hat{V}_{A}U(t_{1})|\psi_{0}\rangle-\frac{1}{2}e^{i\alpha}\hat{V}_{B}\hat{U}(t_{2})|\psi_{0}\rangle\right]\,.\end{split}$ (17)
@@
$2\left|\left\langle 0\middle|\mathopen{}\Psi_{V_{A}V_{B}}^{\alpha}(t_{1},t_{2})\right\rangle\right|^{2}-1\\ =\frac{1}{2}e^{-i\alpha}\big{\langle}\psi_{0}\,\big{|}\,\mathopen{}\hat{U}^{\dagger}(t_{2})\hat{V}^{\dagger}_{B}\hat{U}(t_{2}-t_{1})\hat{V}_{A}\hat{U}(t_{1})\,\big{|}\,\mathopen{}\psi_{0}\big{\rangle}+\text{c.c.}\,,$ (18)
@@
where c.c. denotes the complex conjugate. By setting $\alpha=\alpha_{+}\coloneqq{}0$ and $\alpha=\alpha_{-}\coloneqq{}\pi/2$, we get the real and the imaginary parts of the dynamical correlation function of $\hat{V}_{A}$ and $\hat{V}^{\dagger}_{B}$, respectively:
$1\displaystyle 2P_{V_{A}V_{B}}^{+}(t_{1},t_{2})-$ 1 $=\frac{1}{2}\big{\langle}\hat{V}_{A}^{\dagger}(t_{1})\hat{V}_{B}(t_{2})+\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{2})\big{\rangle}$
$=\big{\langle}\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{1})\big{\rangle}\,,$ (19a)
$1\displaystyle 2P_{V_{A}V_{B}}^{-}(t_{1},t_{2})-$ 1 $=\frac{1}{2}i\big{\langle}\hat{V}_{A}^{\dagger}(t_{1})\hat{V}_{B}(t_{2})-\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{1})\big{\rangle}$
$=\operatorname{Im}\big{\langle}\hat{V}_{B}^{\dagger}(t_{2})\hat{V}_{A}(t_{1})\big{\rangle}\,.$ (19b)
By means of the decomposition in Eq. 5, we can express the two-time (anti-)commutator as
$C_{AB}^{\pm}$ $=\frac{1}{4}\lVert\hat{A}\rVert\lVert\hat{B}\rVert\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\big{\langle}e^{i\alpha_{\pm}}\hat{V}_{A}\hat{V}_{B}^{\prime}+e^{-i\alpha_{\pm}}\hat{V}_{B}^{\prime}\hat{V}_{A}\big{\rangle}$
$=\frac{1}{4}\lVert\hat{A}\rVert\lVert\hat{B}\rVert\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\big{\langle}e^{-i\alpha_{\pm}}\hat{V}_{B}^{\prime\,\dagger}\hat{V}_{A}+e^{i\alpha_{\pm}}\hat{V}_{A}^{\dagger}\hat{V}_{B}^{\prime}\big{\rangle}$
$=\frac{1}{2}\lVert\hat{A}\rVert\lVert\hat{B}\rVert\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\big{(}\langle\hat{V}_{B}^{\prime\,\dagger}\hat{V}_{A}\rangle\cos\alpha_{\pm}$
[…]
where in the second line we have rearranged the terms in the sum. Combining this expression with Eq. 19 finally yields the desired result in Eq. 6.
Appendix B Variance of the Hadamard test